OpenStudy (anonymous):

Which will give me the incorrect slope, y2-y1/x2-x1, y2-y1/x1-x2, y1-y2/x1-x2, rise/run?

hartnn (hartnn):

\(\huge m=\frac{y_1-y_2}{x_1-x_2} = \frac{y_2-y_1}{x_2-x_1} \)